# Counting Equilibria of the Electrostatic Potential

### v1

@@ Abstract.
A century and a half ago, James C. Maxwell [8] conjectured that the number of zeroes of the electric field (equilibria of the potential) generated by a collection of $n$ point charges is at most $(n-1)^{2}$. In 2007 Gabrielov, Novikov and Shapiro proved a much larger upper bound and posed two further conjectures [7]. It is quite possible that Maxwell’s quadratic upper bound is not tight, so it is prudent to find lower bounds. Our main contributions in this article are:
  • the construction of examples that achieve the highest ratios of number of electric field zeroes by point charges found to this day;
  • a counterexample to Conjecture 1.8 in [7] that the number of equilibria cannot exceed those of the distance function defined by the unit point charges.
@@ 1. Introduction
Context. Let $A_{1},A_{2},\ldots,A_{n}$ be points in $\mathbb{R}^{3}$ with corresponding real charges $\zeta_{1},\zeta_{2},\ldots,\zeta_{n}$. We call $A_{i}$ together with $\zeta_{i}$ a point charge. The electrostatic potential generated by the $n$ point charges is the function $V\colon\mathbb{R}^{3}\setminus\{A_{1},A_{2},\ldots,A_{n}\}\to\mathbb{R}$ defined by
@@
A general upper bound on the number of equilibria was only obtained in 2013 [7]. However, it grows much faster than the quadratic upper bound conjectured by Maxwell. Indeed, already for $n=3$ this bound is $139314069504$. Beyond the upper bound, Gabrielov, Novikov and Shapiro prove interesting results and pose stimulating conjectures, which we will now explain and later address in this article. An important auxiliary tool in [7] is the family of potentials
@@
Main Results. As stated in the abstract, the key contributions of this article are a new record for the ratio of the number of equilibria over the number of point charges, a counterexample to Conjecture 2, and a number of other interesting results and explorations. In Sections 2, we give a general discussion and exploration of examples obtained by placing unit charges at the vertices of several polyhedra. It is here that we arrive at our first main result, which is a family of examples constructed in Sections 2.3 and 2.4. The conclusion of that construction is the following.
Result 1 (New record for number of equilibria).
@@
Result 2 (Counterexample to Conjecture 2).
@@ 2.3. Prisms and Anti-prism
For $k=6$, this is the same ratio we get for the cuboctahedron after perturbing the point charges so that the degenerate equilibrium at the center unfolds into individual non-degenerate equilibria; see Section 2.4 for details. To maximize this ratio, we minimize the number of sides, which suggests that the $4$-sided or square anti-prism is our best choice. In contrast, the equilibria for the $3$-sided or triangular anti-prism are either too close to distinguish visually or are absorbed by a degenerate equilibrium at the center.
@@
In addition to the results on the maximum number of equilibria for anti-prims, we discover an interesting universal phenomenon: the existence of a transition that occurs when the relative height of the prism or anti-prism is $\sqrt{2}$. This transition is universal in the sense that it occurs for both prisms and anti-prims formed with any regular $k$-gon and always at relative height $\sqrt{2}$. We prove this in Appendix C and describe it here together with a result currently limited to the special cases of square and a triangular anti-prisms.
@@
    Furthermore assuming a square or triangular anti-prism and a relative height smaller than $\sqrt{2}$, there is an equilibrium on every line segment connecting the center of the anti-prism to the midpoint of a lateral edge.
@@ 2.4. Increasing the Ratio
In particular, the ratio of the number of equilibria by electric charges in the $\ell$-th iteration is given by $\frac{m}{n-1}\cdot\frac{n_{\ell}-1}{n_{\ell}}$, which for $\ell=1$ coincides with the initial ratio of $\frac{m}{n}$, and as $\ell\to+\infty$ converges to $\frac{m}{n-1}$. In all examples we have computed, the iterated square anti-prism achieves the largest fraction $\nicefrac{{25}}{{7}}$ in the limit, when $\ell\to\infty$. This is the highest ratio ever observed by far. In particular, we find that for any $\varepsilon>0$, there is an iteration $\ell_{*}$ such that all further iterations $\ell>\ell_{*}$ have a ratio greater than $\nicefrac{{25}}{{7}}-\varepsilon$. This establishes Result 1 and naturally leads to the question of whether one can find an example of a configuration which achieves a greater ratio. We state this inquiry as follows.
@@ 3.1. No Maxima
Again, the proof of this result will be presented in Appendix D. Finally, we point out that in this and other examples, the origin becomes a non-degenerate equilibrium for $p>1$. Thus, in terms of non-degeneracy, increasing the power, $p$, can have a similar effect as perturbing the point charges; see [11, Thm 6.3].
@@ 3.2. Counterexample
The hope expressed as Conjecture 1.8 in [7]—and stated as Conjecture 2 in this article—is that $V\colon\mathbb{R}^{3}\to\mathbb{R}$ cannot have more equilibria than $E\colon\mathbb{R}^{3}\to\mathbb{R}$, in which the latter are defined using the limit process in (10). As stated in Result 2, the unit point charges at the vertices of the truncated octahedron form a counterexample to this conjecture, and we shall now explain why. This Voronoi domain is also known as the truncated octahedron; see Figure 4, which also shows the eqilibria of the electrostatic potential defined by placing unit point charges at its vertices. This polytope has $14$ faces ($6$ squares and $8$ hexagons) and $36$ edges ($24$ shared by a square and a hexagon and $12$ shared by two hexagons). Correspondingly, $E$ has $14$ $1$-saddles and $36$ $2$-saddles, which we compare to the $18$ $1$-saddles and $36$ $2$-saddles of $V$. Hence, the Euclidean distance function has fewer $1$-saddles than the electrostatic potential of the same unit point charges, and it has equally many $2$-saddles. In total, $V$ has more equilibria than $E$, which contradicts Conjecture 1.8 in [7].

### v2

@@ Abstract.
In 1873, James C. Maxwell [10] conjectured that the electric field generated by $n$ point charges in generic position has at most $(n-1)^{2}$ isolated zeroes.
The first (non-optimal) upper bound was only obtained in 2007 by Gabrielov, Novikov and Shapiro, who also posed two additional interesting conjectures [7].
In this article we give the best upper bound known to date on the number of zeroes of the electric field, and construct a counterexample to Conjecture 1.8 in [7] that the number of equilibria cannot exceed those of the distance function defined by the unit point charges.
Finally, we note that it is quite possible that Maxwell’s quadratic upper bound is not tight, so it is prudent to find lower bounds. Hence, we also explore examples and construct configurations of charges achieving the highest ratios of the number of electric field zeroes by point charges found to this day.
@@ 1. Introduction
Context. The electrostatic potential generated by $n$ point charges located at $A_{1},A_{2},\ldots,A_{n}\in\mathbb{R}^{3}$ of magnitudes $\zeta_{1},\zeta_{2},\ldots,\zeta_{n}\in\mathbb{R}$ is the function $V\colon\mathbb{R}^{3}\setminus\{A_{1},A_{2},\ldots,A_{n}\}\to\mathbb{R}$ defined by
@@
A general upper bound on the number of equilibria was only obtained in 2007 [7], by using Kovanskii’s theory of Fewnomials [9]. However, it grows much faster than the quadratic upper bound conjectured by Maxwell. Indeed, it is given by $2^{2n^{2}}(3n)^{2n}$, which already for $n=3$ is $139314069504$. More recently, using the Thom-Milnor theorem, this was further improved in [18] to $5\times 9^{3+n}$. As we shall see further below, one of our contributions in this article will be to improve this bound even further by a clever use of Bezout’s theorem.
Beyond the upper bound, in [7] Gabrielov, Novikov and Shapiro prove interesting results and pose very stimulating conjectures, which we will now explain and later address in this article. An important auxiliary tool in [7] is the family of potentials
@@
Main Results. As stated in the abstract, the key contributions of this article are:
  • an improved upper bound on the number of isolated equilibria for the electrostatic potential;
  • a new record for the ratio of the number of equilibria over the number of point charges;
  • a counterexample to Conjecture 2; and
  • a number of other interesting results and explorations.
We shall now start by stating our first main result, namely the new upper bound on the number of isolated equilibria for the electrostatic potential.
Result 1.
Let $n\in\mathbb{N}$, $A_{2},\ldots,A_{n}\in\mathbb{R}^{3}$, and $\0\zeta_{1},\ldots,\zeta_{n}\in\mathbb{R}\backslash\{0\}$. Then, for the generic position of $A_{1}\in\mathbb{R}^{3}$, the number of isolated critical points of $V$ is at most
$2^{n}\times(3n-2)^{3}.$
The proof of this result appears in Section 2, together with an analogous but stronger result for the potentials $V_{p}$ for even $p$. For these, an easier application of the same method yields a polynomial upper bound.
In Sections 3, we give a general discussion and exploration of examples obtained by placing unit charges at the vertices of several polyhedra. It is here that we arrive at our first main result, which is a family of examples constructed in Sections 3.3 and 3.4. The conclusion of that construction is the following.
Result 2 (New record for number of equilibria).
@@
Result 3 (Counterexample to Conjecture 2).
@@
2. The upper bound
This section contains a proof of the upper bound stated in our first main result 1 is as follows. We start by computing the critical point equation and clearing out denominators in Lemma 1. The resulting equations can be turned into polynomial ones by a trick that formally introduces extra variables to encode the quantities given by square roots. This is the content of section 2.1, where the key polynomial maps are defined.
To give an upper bound on the number of isolated solutions to these polynomial equations, we use Bezout’s theorem, which gives such a bound from the degrees of the relevant polynomial equations. However, there is a subtlety involved in using this upper bound. It requires showing that non-degenerate critical points of $V$ correspond to non-isolated zeroes of the polynomial system. This is proven in 2.3 and these preliminary arguments are then all put together in 2.4 to finalize the proof of Theorem 1.
2.1. Towards polynomial equations
The electric potential generated by charges of magnitude $\zeta_{i}$ located at the points $p_{i}$, for $i=1,\ldots,n$, is
$V(x)=\sum_{i=1}^{n}\frac{\zeta_{i}}{r_{i}},$
where $r_{i}=|x-p_{i}|$. We shall write $p_{i}=(p_{i1},p_{i2},p_{i3})$ and $x=(x_{1},x_{2},x_{3})$.
Lemma 1.
For $j=1,2,3$, let
$R_{j}(x)=\sum_{i=1}^{n}\zeta_{i}(x_{j}-p_{ij})\prod_{\ell\neq i}r_{\ell}^{3}.$
A point $x\in\mathbb{R}^{3}$ is a critical point of $V$ if and only if it is a common zero of $R_{1}$, $R_{2}$, and $R_{3}$.
Proof.
We can compute that $\partial_{x_{j}}r_{i}=\frac{x_{j}-p_{ij}}{r_{i}}$, then $\partial_{x_{j}}r_{i}^{-1}=-\frac{x_{j}-p_{ij}}{r_{i}^{3}}$ and therefore
$\partial_{x_{j}}V(x)=-\sum_{i=1}^{n}\frac{\zeta_{i}}{r_{i}^{3}}(x_{j}-p_{ij}).$
We can clear out denominators as follows
$\partial_{x_{j}}V(x)$ $=-\sum_{i=1}^{n}\zeta_{i}(x_{j}-p_{ij})\times\frac{\prod_{\ell\neq i}r_{\ell}^{3}}{\prod_{m=1}^{n}r_{m}^{3}}$
$=-\frac{\sum_{i=1}^{n}\zeta_{i}(x_{j}-p_{ij})\prod_{\ell\neq i}r_{\ell}^{3}}{\prod_{m=1}^{n}r_{m}^{3}}.$
At this point we introduce the variables $u_{1},\ldots,u_{n}$ and rewrite the system from Lemma 1 as the zeros of the following polynomial functions
$\mathcal{R}_{j}(x,u)$ $=\sum_{i=1}^{n}\zeta_{i}(x_{j}-p_{ij})\prod_{\ell\neq i}u_{\ell}^{3},\ \ \text{for}\ j\in\{1,2,3\},$
$\mathcal{Q}_{m}(x,u)$ $=u_{m}^{2}-|x-p_{m}|^{2},\ \ \text{for}\ m\in\{1,\ldots,n\}.$
Hence, $x$ is a critical point of $V$ if and only if there is $u\in\mathbb{R}^{n}$ such that $(x,u)$ is a zero of the polynomial map
$\mathcal{P}:\mathbb{R}^{3}\times\mathbb{R}^{n}\to\mathbb{R}^{3}\times\mathbb{R}^{n}$
given by
(5) $\mathcal{P}(x,u)=\left(\mathcal{R}(x,u),\mathcal{Q}(x,u)\right).$
As polynomials in the variables $(x,u)\in\mathbb{R}^{3}\times\mathbb{R}^{n}$, its components have degrees
$\deg(\mathcal{R}_{j})=1+3(n-1)=3n-2,$
for $j\in\{1,2,3\}$, and
$\deg(\mathcal{Q}_{m})=2,$
for $m\in\{1,\ldots,n\}$.
2.2. Bezout’s theorem on affine hypersurfaces
The main tool we shall use is the following theorem.
Theorem 1 (Affine version of Bezout’s theorem).
Let $P_{1},\ldots,P_{m}\in\mathbb{R}[x_{1},\ldots,x_{m}]$ be polynomials of degree $d_{1},\ldots,d_{d}$ respectively. Then, the number of non-degenerate zeroes of the system $P_{1}=0,\ldots,P_{m}=0$ is at most $d_{1}\ldots d_{m}$.
A proof of this result for complex polynomials is given in [17] which states the result for complex polynomials, though the upper bound can be obtained by extending the polynomials to become complex (a non-degenerate zero of a real polynomial equation is also non-degenerate as a zero of its complexification). Alternatively, this result, as stated, is a particular case of the Theorem stated in page 12 of [9]. This is the result used in [7] to give the previously found bound. However, here we are able to substantially improve it by working with polynomial, rather quasi-polynomial, equations.
2.3. Isolated zeroes
The affine Bezout’s theorem will be used to give an upper bound on the number of isolated zeroes of the polynomial map $\mathcal{P}$ defined in 5. However, for this to give an effective bound on the number of non-degenerate zeroes of the electric potential $V$ we must show that they correspond to isolated zeroes of this polynomial map $P$. Our next result shows this to be the case if one of the charges charges can be chosen in generic position..
Lemma 2.
For the generic choice of $p_{1}$, any critical point $x$ of $V$ is such that $(x,r)$ is an isolated zero of $\mathcal{P}$ for $r=(r_{1},\ldots,r_{n})$.
Proof.
The result will be proven by showing that, for generic $p_{1}$ and $x$ a critical point of $V$, the derivative of the map $\mathcal{P}$ has full rank, i.e $n+3$. In the standard basis of $\mathbb{R}^{3+n}$, such derivative is represented by the map
$\begin{pmatrix}\left(\frac{\partial\mathcal{R}_{j}}{\partial x_{s}}\right)_{js}&|&\left(\frac{\partial\mathcal{R}_{j}}{\partial u_{q}}\right)_{jq}\\ ----&|&----\\ \left(\frac{\partial\mathcal{Q}_{m}}{\partial x_{s}}\right)_{ms}&|&\left(\frac{\partial\mathcal{Q}_{m}}{\partial u_{q}}\right)_{mq}\end{pmatrix}.$
We shall now compute the partial derivatives, at the point $(x,r)$, as follows
$\frac{\partial\mathcal{R}_{j}}{\partial x_{s}}=\delta_{js}\sum_{i=1}^{n}\zeta_{i}\prod_{\ell\neq i}r_{\ell}^{3},\ \ \ \ \frac{\partial\mathcal{R}_{j}}{\partial u_{q}}=3u_{q}^{2}\sum_{i=1}^{n}\zeta_{i}\prod_{\ell\neq i,q}r_{\ell}^{3},$
and
$\frac{\partial\mathcal{Q}_{m}}{\partial x_{s}}=-2(x_{r}-p_{ms}),\ \ \ \ \frac{\partial\mathcal{Q}_{m}}{\partial u_{q}}=2r_{m}\delta_{mq}.$
Now, we must have that $r_{\ell}\neq 0$ for all $\ell\in\{1,\ldots,n\}$ as at an equilibrium $x$ cannot coincide with one of the charges $p_{1},\ldots,p_{n}$. Therefore $\left(\frac{\partial\mathcal{Q}_{m}}{\partial u_{q}}\right)_{mq}$ has rank $n$ along the zero set of $\mathcal{P}$. On the other hand, $\left(\frac{\partial\mathcal{R}_{j}}{\partial x_{s}}\right)_{js}$ has rank $3$ provided that
$\sum_{i=1}^{n}\zeta_{i}\prod_{\ell\neq i}r_{\ell}^{3}\neq 0.$
We shall now show that this is the case for the generic choice of $p_{1}\in\mathbb{R}^{3}$. For that, it is enough to show that the map $F:\mathbb{R}^{3}\to\mathbb{R}$ given by
$F(p_{1})=\sum_{i=1}^{n}\zeta_{i}\prod_{\ell\neq i}r_{\ell}^{3},$
is a submersion at $p_{1}\in\mathbb{R}^{3}$ such that $F(p_{1})=0$ and $x\in\mathbb{R}^{3}$ as above. If the map $F$ is not a local submersion at such a point, all its partial derivatives must vanish. Then, by computing
$\frac{\partial F}{\partial p_{1s}}$ $=\sum_{i=2}^{n}\zeta_{i}2r_{1}^{2}\frac{\partial r_{1}}{\partial p_{1s}}\prod_{\ell\neq 1,i}r_{\ell}^{3}$
$=-2\sum_{i=2}^{n}\zeta_{i}r_{1}(x_{s}-p_{1s})\prod_{\ell\neq 1,i}r_{\ell}^{3}$
$=-2r_{1}(x_{s}-p_{1s})\sum_{i=2}^{n}\zeta_{i}\prod_{\ell\neq 1,i}r_{\ell}^{3}.$
we find that to be the case if and only if either $p_{1}=x$ or $\sum_{i=2}^{n}\zeta_{i}\prod_{\ell\neq 1,i}r_{\ell}^{3}=0$. As mentioned before, the first condition is impossible as an equilibrium point cannot coincide with one of the charges. As for the second, together with $F(p_{1})=0$, it would imply that both
$\sum_{i=1}^{n}\zeta_{i}\prod_{\ell\neq i}r_{\ell}^{3}$ $=0$
$\sum_{i=2}^{n}\zeta_{i}\prod_{\ell\neq 1,i}r_{\ell}^{3}$ $=0.$
The first equation above can also be written as
$\zeta_{1}\prod_{\ell\neq 1}r_{\ell}^{3}+r_{1}^{3}\sum_{i=2}^{n}\zeta_{i}\prod_{\ell\neq 1,i}r_{\ell}^{3}=0.$
Then, subtracting to it $r_{1}^{3}$ times the second equation gives
$\zeta_{1}\prod_{\ell\neq 1}r_{\ell}^{3}=0,$
which again, is impossible because an equilibrium point $x$ can not coincide with the location of a charge. ∎
2.4. Finalizing the proof of Result 1
By construction, and due to Lemma 2, the zeros of $\nabla V$ are in correspondence to isolated zeroes of
$\mathcal{P}:\mathbb{R}^{3+n}\to\mathbb{R}^{3+n},$
for the generic choice of the point $p_{1}$. This is at most the number of isolated zeroes of its analytic extension $P:\mathbb{C}^{3+n}\to\mathbb{C}^{3+n}$ and Bezout’s theorem 1 guarantees that the number of such zeroes, counted with multiplicity, is the product of the degrees
$\deg(\mathcal{R}_{1})\deg(\mathcal{R}_{2})\deg(\mathcal{R}_{3})\times\deg(\mathcal{Q}_{1})\ldots\deg(\mathcal{Q}_{n}).$
These degrees where computed in subsection 2.1 where we found this to be
$(3n-2)^{3}\times 2^{n}.$
Remark 1.
In [8], Bezout’s theorem is also used to obtain an upper bound on the number of equilibria for planar charges. The procedure is, however very different. First, we consider any type of generically placed point charges, not just planar ones. Secondly, our procedure allows us to explicitly prove Lemma 2, which shows that the non-degenerate equilibria become isolated zeroes of the polynomial map $\mathcal{P}$. Thus allowing us to use Bezout’s theorem to obtain an effective upper bound.
2.5. The modified potentials $V_{p}$
For $p>0$ let us instead consider the potential
$V_{p}(x)=\sum_{i=1}^{n}\frac{\zeta_{i}}{r_{i}^{p}}.$
Proposition 1.
For even $p=2r\in 2\mathbb{N}$, the number of isolated critical points of $V_{p}$ is at most
$\left(r(n-1)+2-r\right)^{3}.$
In particular, the number of isolated critical points of $V_{2}$ is at most $n^{3}$.
Proof.
Differentiating with respect to $x_{j}$ and clearing out denominators, we obtain
$\frac{\partial V_{p}}{\partial x_{j}}$ $=-\sum_{i=1}^{n}\frac{\zeta_{i}}{r_{i}^{p+2}}(x_{j}-p_{ij})$
$=-\frac{1}{\prod_{\ell=1}^{n}r_{\ell}^{p+2}}\sum_{i=1}^{n}\zeta_{i}(x_{j}-p_{ij})\prod_{m\neq i}r_{m}^{p+2}.$
If $p=2r\in 2\mathbb{N}$ is even, then the numerator above defines a polynomial in $x$
$\tilde{P_{j}}(x)=\sum_{i=1}^{n}\zeta_{i}(x_{j}-p_{ij})\prod_{m\neq i}r_{m}^{2r+2}$
of degree
$1+(n-1)(r-1)=r(n-1)+2-r.$
Again, by complexifying it and using the affine version of Bezout’s theorem we conclude that
#$\#\left(\tilde{P_{1}}^{-1}(0)\cap\tilde{P_{2}}^{-1}(0)\cap\tilde{P_{3}}^{-1}(0)\right)\leq\left(r(n-1)+2-r\right)^{3}.$
In particular, for $r=1$ we obtain the upper bound $n^{3}$. ∎
Remark 2.
The bound obtained in Proposition 1 is an improvement of the previously known bound of
$(1+2(1+r)(n-1))\times(1+4(r+1)(n-1))^{2},$
which had been obtained in [18].
@@ 3.3. Prisms and Anti-prism
For $k=6$, this is the same ratio we get for the cuboctahedron after perturbing the point charges so that the degenerate equilibrium at the center unfolds into individual non-degenerate equilibria; see Section 3.4 for details. To maximize this ratio, we minimize the number of sides, which suggests that the $4$-sided or square anti-prism is our best choice. In contrast, the equilibria for the $3$-sided or triangular anti-prism are always too close to distinguish using our methods.
@@
In addition to the results on the maximum number of equilibria for anti-prims, we discover an interesting universal phenomenon: the existence of a transition that occurs when the relative height of the prism or anti-prism is $\sqrt{2}$. This transition is universal in the sense that it occurs for both prisms and anti-prims formed with any regular $k$-gon and always at relative height $\sqrt{2}$. We prove this in Appendix C and describe it here together with an existence result for some other equilibria.
@@
    Furthermore, if the relative height is smaller than $\sqrt{2}$, there is an equilibrium on every line segment connecting the center of the anti-prism to the midpoint of a lateral edge.
@@ 3.4. Increasing the Ratio
In particular, the ratio of the number of equilibria by electric charges in the $\ell$-th iteration is given by $\frac{m}{n-1}\cdot\frac{n_{\ell}-1}{n_{\ell}}$, which for $\ell=1$ coincides with the initial ratio of $\frac{m}{n}$, and as $\ell\to+\infty$ converges to $\frac{m}{n-1}$. In all examples we have computed, the iterated square anti-prism achieves the largest fraction $\nicefrac{{25}}{{7}}$ in the limit, when $\ell\to\infty$. This is the highest ratio ever observed by far. In particular, we find that for any $\varepsilon>0$, there is an iteration $\ell_{*}$ such that all further iterations $\ell>\ell_{*}$ have a ratio greater than $\nicefrac{{25}}{{7}}-\varepsilon$. This establishes Result 2 and naturally leads to the question of whether one can find an example of a configuration which achieves a greater ratio. We state this inquiry as follows.
@@ 4.1. No Maxima
Again, the proof of this result will be presented in Appendix D. Finally, we point out that in this and other examples, the origin becomes a non-degenerate equilibrium for $p>1$. Thus, in terms of non-degeneracy, increasing the power, $p$, can have a similar effect as perturbing the point charges; see [13, Thm 6.3].
@@ 4.2. Counterexample
The hope expressed as Conjecture 1.8 in [7]—and stated as Conjecture 2 in this article—is that $V\colon\mathbb{R}^{3}\to\mathbb{R}$ cannot have more equilibria than $E\colon\mathbb{R}^{3}\to\mathbb{R}$, in which the latter are defined using the limit process in (11). As stated in Result 3, the unit point charges at the vertices of the truncated octahedron form a counterexample to this conjecture, and we shall now explain why. This Voronoi domain is also known as the truncated octahedron; see Figure 4, which also shows the eqilibria of the electrostatic potential defined by placing unit point charges at its vertices. This polytope has $14$ faces ($6$ squares and $8$ hexagons) and $36$ edges ($24$ shared by a square and a hexagon and $12$ shared by two hexagons). Correspondingly, $E$ has $14$ $1$-saddles and $36$ $2$-saddles, which we compare to the $18$ $1$-saddles and $36$ $2$-saddles of $V$. Hence, the Euclidean distance function has fewer $1$-saddles than the electrostatic potential of the same unit point charges, and it has equally many $2$-saddles. In total, $V$ has more equilibria than $E$, which contradicts Conjecture 1.8 in [7].